# On multi-graded Proj schemes

### v3

@@ 4 Examples arising from reductive algebraic groups
In this section we show that a number of varieties of interest in Geometric Representation Theory can be described as Proj schemes associated with certain graded rings in a canonical way (which, in particular, do not require a specific choice of dominant character). We see these examples as a motivation to study the general theory considered in the rest of the paper, but none of the considerations of the present sections will be used in proofs of results of other sections. (In Section 5, these results will only be used as illustrations of certain general statements.)
@@ 4.1 Flag varieties
In this section we fix an algebraically closed field $𝕜\Bbbk$, [^2] It is likely that most of the results in this section can be generalized to reductive group schemes over more general base schemes, but we will not go in this direction. and a connected reductive algebraic group $\mathbf{G}$ over $𝕜\Bbbk$. We fix a Borel subgroup $\mathbf{B}$ and a maximal torus $\mathbf{T}\subset\mathbf{B}$, and denote by $\mathbf{U}$ the unipotent radical of $\mathbf{B}$. We denote by $\mathbb{X}$ the character lattice of $\mathbf{T}$ (a free $\mathbb{Z}$-module of finite rank), and by $\mathfrak{R}\subset\mathbb{X}$ the root system of $(\mathbf{G},\mathbf{T})$, i.e. the subset of nonzero $\mathbf{T}$-weights in the Lie algebra of $\mathbf{G}$. We will denote by $\mathfrak{R}_{+}\subset\mathfrak{R}$ the subset consisting of the opposites of the $\mathbf{T}$-weights in the Lie algebra of $\mathbf{U}$; this forms a positive system. We will also consider the cocharacter lattice $\mathbb{X}^{\vee}$ of $\mathbf{T}$, which identifies with the dual of $\mathbb{X}$, and the system of coroots $\mathfrak{R}^{\vee}\subset\mathbb{X}^{\vee}$. There exists a canonical bijection $\mathfrak{R}\xrightarrow{\sim}\mathfrak{R}^{\vee}$, which we will as usual denote $\alpha\mapsto\alpha^{\vee}$. Our choice of positive system $\mathfrak{R}_{+}$ lets us define a submonoid $\mathbb{X}_{+}\subset\mathbb{X}$ of dominant weights as
@@
$\Gamma(\mathbf{G}/\mathbf{B},{\mathcal{O}}_{\mathbf{G}/\mathbf{B}}(\lambda))$
vanishes unless $\lambda\in\mathbb{X}^{+}$, and is finite-dimensional in this case; see e.g. [Ja03, §II.2.1 and Proposition II.2.6]. For $\lambda,\mu\in\mathbb{X}$ there exists a canonical isomorphism
@@
$A:=\bigoplus_{\lambda\in\mathbb{X}_{+}}\Gamma(\mathbf{G}/\mathbf{B},{\mathcal{O}}_{\mathbf{G}/\mathbf{B}}(\lambda)).$
Its spectrum $\mathbf{X}:={\operatorname{Spec}}(A)$ is the “affine completion” of the space $\mathbf{G}/\mathbf{U}$; it is a classical object, studied in particular in [AR, §6.2.1] [^3] In this section of [AR] it is assumed that the base field has characteristic 00. This assumption is however not used for the statements we use below. (where it is denoted $\mathcal{X}$).
@@
    A similar description of $\mathbf{G}/\mathbf{B}$ as the Proj-scheme associated with an $\mathbb{N}$-graded ring, which requires to choose a strictly dominant weight, is classical, although we were not able to find an explicit mention in the published literature. For an exposition of the proof, see the notes [Wa].
@@
We will give two independent proofs of Proposition 4.1. The first one uses the description of the Proj-scheme as a geometric quotient under the action of a diagonalizable group scheme explained in §3.1. The second one uses results of [BS03] on ample families of line bundles.
@@
It is clear that $\bigoplus_{\lambda\in\mathbb{X}_{++}}\Gamma(\mathbf{G}/\mathbf{B},{\mathcal{O}}_{\mathbf{G}/\mathbf{B}}(\lambda))$ is an ideal in $A$. If $f\in A$ is a homogenous relevant element, then $\deg(f)$ must be strictly dominant; in fact, if there exists $\alpha\in\mathfrak{R}_{+}$ such that $\langle\deg(f),\alpha^{\vee}\rangle=0$, then any homogeneous divisor $g$ of a power of $f$ will also satisfy $\langle\deg(g),\alpha^{\vee}\rangle=0$, contradicting the fact that $f$ is relevant. As a consequence, we have
@@
Any $\lambda\in\mathbb{X}_{0}$ extends to a character of $\mathbf{G}$, so that the space $\Gamma(\mathbf{G}/\mathbf{B},{\mathcal{O}}_{\mathbf{G}/\mathbf{B}}(\lambda))$ is $1$-dimensional (by the tensor identity [Ja03, Proposition I.3.6]), and its nonzero elements are invertible.
@@
By [BS03, Proposition 4.2], there exists a canonical rational map from $\mathbf{G}/\mathbf{B}$ to ${\operatorname{Proj}}^{\mathbb{X}}(A)$; what we will show is that this rational morphism is an isomorphism.
@@
In this case, we have described the ideal $A_{\dagger}\subset A$ in Lemma 4.3. This ideal coincides with that considered in [AR, Remark 6.2.3]; comparing these description we obtain an identification
@@
Decomposing the space ${\mathcal{O}}(\mathbf{G}/\mathbf{U})$ according to the weight spaces of the natural action of $\mathbf{T}$ and looking at the definitions we obtain an identification
${\mathcal{O}}(\mathbf{G}/\mathbf{U})=\bigoplus_{\lambda\in\mathbb{X}}\Gamma(\mathbf{G}/\mathbf{B},{\mathcal{O}}_{\mathbf{G}/\mathbf{B}}(\lambda))=A.$
Hence we have ${\operatorname{Spec}}(A)=(\mathbf{G}/\mathbf{U})_{\mathrm{aff}}$. The above proof amounts to noticing that, in case $\mathbf{G}$ has simply connected derived subgroup, the natural morphism
$\mathbf{G}/\mathbf{U}\to(\mathbf{G}/\mathbf{U})_{\mathrm{aff}}$
is an open immersion, with image ${\operatorname{Spec}}(A)\smallsetminus V(A_{\dagger})$.
@@
${\operatorname{Proj}}^{\mathbb{X}}(A)\hookrightarrow{\operatorname{Proj}}^{\mathbb{X}}\left(\bigotimes_{\alpha\in\mathfrak{R}_{\mathrm{s}}}\mathrm{S}(A_{\varpi_{\alpha}})\right)$
@@
Fix such $\lambda$. If $a\in A_{\lambda}$, there exist elements $(b_{\sigma}:\sigma\in E)$ in $\widetilde{A}_{\lambda-\sum_{\alpha}\varpi_{\alpha}}$ such that $a=\sum_{\sigma}b_{\sigma}\cdot f_{\sigma}$. Then if $N>|E|\cdot\max_{\sigma}n_{\sigma}$, we have $a^{N}=\sum_{\sigma}c_{\sigma}g_{\sigma}$ for some homogeneous elements $c_{\sigma}\in\widetilde{A}$, where $\deg(c_{\sigma})+\deg(g_{\sigma})=N\lambda$ for any $\sigma$. Here we have $\deg(c_{\sigma})\in\mathbb{X}$, hence $c_{\sigma}\in A$, and the desired claim follows.
@@ 4.2 Some vector bundles over the flag variety
i.e. the quotient of the product $\mathbf{G}\times V$ by the (free) action of $\mathbf{B}$ defined by $b\cdot(g,x)=(gb^{-1},b\cdot x)$. (This construction is a special case of that considered in [Ja03, §II.5.14].) It is a vector bundle over $\mathbf{G}/\mathbf{B}$ (in particular, a smooth variety). For $\lambda\in\mathbb{X}$ we will denote by ${\mathcal{O}}_{\mathbf{G}\times^{\mathbf{B}}V}(\lambda)$ the pullback to $\mathbf{G}\times^{\mathbf{B}}V$ of the line bundle ${\mathcal{O}}_{\mathbf{G}/\mathbf{B}}(\lambda)$. By (4.1), for $\lambda,\mu\in\mathbb{X}$ we have a canonical isomorphism
@@
and the right-hand side identifies with $\mathbf{G}/\mathbf{B}\times\widetilde{V}$ (via the map $[g,x]\mapsto(g\mathbf{B},g\cdot x)$ for $g\in G$ and $x\in\widetilde{V}$). For any $\lambda\in\mathbb{X}_{+}$ we deduce a canonical morphism
@@
In case $V=\widetilde{V}=\{0\}$, Proposition 4.7 recovers Proposition 4.1 (but our proof will use the latter statement). Another interesting case is when $\widetilde{V}$ is the Lie algebra $\mathbf{g}$ of $\mathbf{G}$ and $V$ is the Lie algebra $\mathbf{u}$ of the unipotent radical of $\mathbf{B}$. In this case, $\mathbf{G}\times^{\mathbf{B}}V$ is the so-called Springer resolution.
@@
There exists $N\in\mathbb{Z}_{>0}$ such that for any $\lambda\in\mathbb{X}^{+}$ which satisfies $\langle\lambda,\alpha^{\vee}\rangle\geqslant N$ for any simple root $\alpha$, the morphism (4.10) is surjective.
@@
Now $\mathcal{V}^{\bot}$ is the vector bundle on $\mathbf{G}/\mathbf{B}$ associated with the orthogonal $V^{\bot}\subset\widetilde{V}^{*}$ of $V\subset\widetilde{V}$. Hence, if we denote by $\Lambda\subset\mathbb{X}$ the (finite) subset consisting of the $T$-weights appearing in the various exterior powers of $V^{\bot}$, then each $\wedge^{i}\mathcal{V}^{\bot}$ is an iterated extension of line bundles of the form ${\mathcal{O}}_{\mathbf{G}/\mathbf{B}}(\mu)$ with $\mu\in\Lambda$. If we choose $N$ such that $\langle\mu,\alpha^{\vee}\rangle\geqslant-N$ for any $\mu\in\Lambda$ and any simple root $\alpha$, then if $\lambda\in\mathbb{X}_{+}$ satisfies $\langle\lambda,\alpha^{\vee}\rangle\geqslant N$ for any simple root $\alpha$, each weight $\lambda+\mu$ with $\mu\in\Lambda$ is dominant. Since line bundles on $\mathbf{G}/\mathbf{B}$ associated with dominant weights have no higher cohomology (this is Kempf’s vanishing theorem, see [Ja03, Proposition II.4.5]), for such $\lambda$ we have
@@
In fact, as in the proof of Lemma 4.3 one sees that the degree of any relevant element $f$ in $A_{V}$ belongs to $\mathbb{X}_{++}$. If $N$ satisfies the condition in Lemma 4.9, then $f^{N}$ belongs to the image of (4.10) for $\lambda=N\deg(f)$, hence to the ideal generated by the elements $\tilde{f}_{\sigma}$. This implies the claim in view of Corollary 3.10.
@@
provided by Proposition 3.12, where the identification is provided by Corollary 3.17. By (4.13) this morphism is defined everywhere. We claim that it is a closed immersion. In fact, by Remarks 3.18 and 4.6 we have
@@
Therefore, to conclude it suffices to prove that the morphism ${\mathcal{O}}(\widetilde{V})\otimes A_{(f_{\sigma})}\to(A_{V})_{(\tilde{f}_{\sigma})}$ is surjective. But, if $N$ is as above any element in $(A_{V})_{(\tilde{f}_{\sigma})}$ can be written in the form $\frac{a}{(\tilde{f}_{\sigma})^{kN}}$ for some $k>0$, where $a\in A$ is homogeneous of degree $kN\deg(\tilde{f}_{\sigma})$. Then $a$ belongs to the image of ${\mathcal{O}}(\widetilde{V})\otimes A$, which implies the desired surjectivity.

### v4

@@ 4 Examples arising from reductive algebraic groups
In this section we show that a number of varieties of interest in Geometric Representation Theory can be described as Proj schemes associated with certain graded rings in a canonical way (which, in particular, does not require a specific choice of dominant character). We see these examples as a motivation to study the general theory considered in the rest of the paper, but none of the considerations of the present sections will be used in proofs of results of other sections. (In Section 5, these results will only be used as illustrations of certain general statements.)
@@ 4.1 Flag varieties
In this section we fix an algebraically closed field $𝕜\Bbbk$, [^4] It is likely that most of the results in this section can be generalized to reductive group schemes over more general base schemes, but we will not go in this direction. and a connected reductive algebraic group $\mathbf{G}$ over $𝕜\Bbbk$. We fix a Borel subgroup $\mathbf{B}\subset\mathbf{G}$ and a maximal torus $\mathbf{T}\subset\mathbf{B}$, and denote by $\mathbf{U}$ the unipotent radical of $\mathbf{B}$. We denote by $\mathbb{X}$ the character lattice of $\mathbf{T}$ (a free $\mathbb{Z}$-module of finite rank), and by $\mathfrak{R}\subset\mathbb{X}$ the root system of $(\mathbf{G},\mathbf{T})$, i.e. the subset of nonzero $\mathbf{T}$-weights in the Lie algebra of $\mathbf{G}$. We will denote by $\mathfrak{R}_{+}\subset\mathfrak{R}$ the subset consisting of the opposites of the $\mathbf{T}$-weights in the Lie algebra of $\mathbf{U}$; this forms a positive system. We will also consider the cocharacter lattice $\mathbb{X}^{\vee}$ of $\mathbf{T}$, which identifies with the $\mathbb{Z}$-dual of $\mathbb{X}$, and the system of coroots $\mathfrak{R}^{\vee}\subset\mathbb{X}^{\vee}$. There exists a canonical bijection $\mathfrak{R}\xrightarrow{\sim}\mathfrak{R}^{\vee}$, which we will as usual denote $\alpha\mapsto\alpha^{\vee}$. Our choice of positive system $\mathfrak{R}_{+}$ lets us define a submonoid $\mathbb{X}_{+}\subset\mathbb{X}$ of dominant weights as
@@
$A:=\bigoplus_{\lambda\in\mathbb{X}_{+}}\Gamma(\mathbf{G}/\mathbf{B},{\mathcal{O}}_{\mathbf{G}/\mathbf{B}}(\lambda)),$
where multiplication is provided by the morphisms (4.2). Its spectrum $\mathbf{X}:={\operatorname{Spec}}(A)$ is the “affine completion” of the space $\mathbf{G}/\mathbf{U}$; it is a classical object, whose main properties are summarized in [AR, §6.2.1] [^5] In this section of [AR] it is assumed that the base field has characteristic 00. This assumption is however not used for the statements we use below. (where it is denoted $\mathcal{X}$).
Lemma 4.1.
The ring $A$ is a domain.
Proof.
Consider the quotient $\mathbf{G}/\mathbf{U}$. Decomposing the space ${\mathcal{O}}(\mathbf{G}/\mathbf{U})$ according to the weight spaces of the natural action of $\mathbf{T}$ and looking at the definitions we obtain an identification
${\mathcal{O}}(\mathbf{G}/\mathbf{U})=\bigoplus_{\lambda\in\mathbb{X}}\Gamma(\mathbf{G}/\mathbf{B},{\mathcal{O}}_{\mathbf{G}/\mathbf{B}}(\lambda))=A.$ (4.3)
This implies the claim, since $\mathbf{G}/\mathbf{U}$ is a smooth variety. ∎
@@
    A similar description of $\mathbf{G}/\mathbf{B}$ as the Proj-scheme associated with an $\mathbb{N}$-graded ring, which requires a choice of a strictly dominant weight, is classical, although we were not able to find an explicit mention in the published literature. For an exposition of the proof, see the notes [Wa].
@@
We will give two independent proofs of Proposition 4.2. The first one uses the description of the Proj-scheme as a geometric quotient under the action of a diagonalizable group scheme recalled in §3.2. The second one uses results of [BS03] on ample families of line bundles.
@@
It is clear that $\bigoplus_{\lambda\in\mathbb{X}_{++}}\Gamma(\mathbf{G}/\mathbf{B},{\mathcal{O}}_{\mathbf{G}/\mathbf{B}}(\lambda))$ is an ideal in $A$; so to prove the inclusion it suffices to prove that the degree of any nonzero relevant element in $A$ belongs to $\mathbb{X}_{++}$. In fact if $f$ is nonzero and homogeneous and if there exists $\alpha\in\mathfrak{R}_{+}$ such that $\langle\deg(f),\alpha^{\vee}\rangle=0$, then any homogeneous divisor $g$ of a power of $f$ (which is necessarily nonzero by Lemma 4.1) will also satisfy $\langle\deg(g),\alpha^{\vee}\rangle=0$, so that $f$ is not relevant.
@@
Any $\lambda\in\mathbb{X}_{0}$ extends to a character of $\mathbf{G}$, so that the space $\Gamma(\mathbf{G}/\mathbf{B},{\mathcal{O}}_{\mathbf{G}/\mathbf{B}}(\lambda))$ is $1$-dimensional (by the tensor identity [Ja03, Proposition I.3.6]), and its nonzero elements are invertible in $A$.
@@
By [BS03, Proposition 4.2], there exists a canonical rational map from $\mathbf{G}/\mathbf{B}$ to ${\operatorname{Proj}}^{\mathbb{X}}(A)$; what we will show is that this rational morphism is defined everywhere, and an isomorphism.
@@
In this case, we have described the ideal $A_{\dagger}\subset A$ in Lemma 4.4. This ideal coincides with that considered in [AR, Remark 6.2.3]; in view of this statement we deduce an identification
@@
In view of the isomorphism (4.3), the above proof amounts to noticing that, in case $\mathbf{G}$ has simply connected derived subgroup, the natural morphism $\mathbf{G}/\mathbf{U}\to(\mathbf{G}/\mathbf{U})_{\mathrm{aff}}$ is an open immersion, with image ${\operatorname{Spec}}(A)\smallsetminus V(A_{\dagger})$.
@@
${\operatorname{Proj}}^{\mathbb{X}}(A)\hookrightarrow{\operatorname{Proj}}^{\mathbb{X}}\left(\bigotimes_{\alpha\in\mathfrak{R}_{\mathrm{s}}}\mathrm{S}(A_{\varpi_{\alpha}})\right),$
@@
  1. 1.
    In the proof above, the fact that ${\operatorname{Proj}}^{\mathbb{X}}(A)$ is separated can also be deduced from Proposition 3.7.
  2. 2.
@@
Fix such $\lambda$. If $a\in A_{\lambda}$, there exist elements $(b_{\sigma}:\sigma\in E)$ in $\widetilde{A}_{\lambda-\sum_{\alpha}\varpi_{\alpha}}$ such that $a=\sum_{\sigma}b_{\sigma}\cdot f_{\sigma}$. Then if $N>|E|\cdot(\max_{\sigma}n_{\sigma}-1)$, we have $a^{N}=\sum_{\sigma}c_{\sigma}g_{\sigma}$ for some homogeneous elements $c_{\sigma}\in\widetilde{A}$, where $\deg(c_{\sigma})+\deg(g_{\sigma})=N\lambda$ for any $\sigma$ such that $c_{\sigma}g_{\sigma}\neq 0$. Here we have $\deg(c_{\sigma})\in\mathbb{X}$, hence $c_{\sigma}\in A$, and the desired claim follows.
@@ 4.2 Some vector bundles over the flag variety
i.e. the quotient of the product $\mathbf{G}\times V$ by the (free) action of $\mathbf{B}$ defined by $b\cdot(g,x)=(gb^{-1},b\cdot x)$. (This construction is a special case of that considered in [Ja03, §I.5.14].) It is a vector bundle over $\mathbf{G}/\mathbf{B}$ (in particular, a smooth variety). For $\lambda\in\mathbb{X}$ we will denote by ${\mathcal{O}}_{\mathbf{G}\times^{\mathbf{B}}V}(\lambda)$ the pullback to $\mathbf{G}\times^{\mathbf{B}}V$ of the line bundle ${\mathcal{O}}_{\mathbf{G}/\mathbf{B}}(\lambda)$. By (4.1), for $\lambda,\mu\in\mathbb{X}$ we have a canonical isomorphism
@@
Remark 4.8.
One can also consider the quotient $\mathbf{G}\times^{\mathbf{U}}V$ of the restriction of the action above to $\mathbf{U}$, which is a (Zariski locally trivial) $\mathbf{T}$-torsor over $\mathbf{G}\times^{\mathbf{B}}V$. As in the proof of Lemma 4.1 one can see that $A_{V}$ identifies with a subring of $\mathcal{O}(\mathbf{G}\times^{\mathbf{U}}V)$, hence is a domain.
@@
and the right-hand side identifies with $\mathbf{G}/\mathbf{B}\times\widetilde{V}$ (via the map $[g,x]\mapsto(g\mathbf{B},g\cdot x)$ for $g\in\mathbf{G}$ and $x\in\widetilde{V}$). For any $\lambda\in\mathbb{X}_{+}$ we deduce a canonical morphism
@@
In case $V=\widetilde{V}=\{0\}$, Proposition 4.9 recovers Proposition 4.2 (but our proof will use the latter statement). Another interesting case is when $\widetilde{V}$ is the Lie algebra $\mathbf{g}$ of $\mathbf{G}$ and $V$ is the Lie algebra $\mathbf{u}$ of the unipotent radical of $\mathbf{B}$. In this case, $\mathbf{G}\times^{\mathbf{B}}V$ is the so-called Springer resolution; see §4.3 below for more details about this case.
@@
There exists $N\in\mathbb{Z}_{>0}$ such that for any $\lambda\in\mathbb{X}_{+}$ which satisfies $\langle\lambda,\alpha^{\vee}\rangle\geqslant N$ for any simple root $\alpha$, the morphism (4.11) is surjective.
@@
Now $\mathcal{V}^{\bot}$ is the vector bundle on $\mathbf{G}/\mathbf{B}$ associated with the orthogonal $V^{\bot}\subset\widetilde{V}^{*}$ of $V\subset\widetilde{V}$. Hence, if we denote by $\Lambda\subset\mathbb{X}$ the (finite) subset consisting of the $\mathbf{T}$-weights appearing in the various exterior powers of $V^{\bot}$, then each $\wedge^{i}\mathcal{V}^{\bot}$ is an iterated extension of line bundles of the form ${\mathcal{O}}_{\mathbf{G}/\mathbf{B}}(\mu)$ with $\mu\in\Lambda$. If we choose $N$ such that $\langle\mu,\alpha^{\vee}\rangle\geqslant-N$ for any $\mu\in\Lambda$ and any simple root $\alpha$, then if $\lambda\in\mathbb{X}_{+}$ satisfies $\langle\lambda,\alpha^{\vee}\rangle\geqslant N$ for any simple root $\alpha$, each weight $\lambda+\mu$ with $\mu\in\Lambda$ is dominant. Since line bundles on $\mathbf{G}/\mathbf{B}$ associated with dominant weights have no higher cohomology (this is Kempf’s vanishing theorem, see [Ja03, Proposition II.4.5]), for such $\lambda$ we have
@@
In fact, as in the proof of Lemma 4.4 one sees that the degree of any nonzero relevant element $f$ in $A_{V}$ belongs to $\mathbb{X}_{++}$. If $N$ satisfies the condition in Lemma 4.11, then $f^{N}$ belongs to the image of (4.11) for $\lambda=N\deg(f)$, hence to the ideal generated by the elements $\tilde{f}_{\sigma}$. This implies the claim in view of Corollary 3.10.
@@
provided by Proposition 3.12, where the identification is provided by Corollary 3.17. By (4.14) this morphism is defined everywhere. We claim that it is a closed immersion. In fact, by Remarks 3.18 and 4.7 we have
@@
Therefore, to conclude it suffices to prove that the morphism ${\mathcal{O}}(\widetilde{V})\otimes A_{(f_{\sigma})}\to(A_{V})_{(\tilde{f}_{\sigma})}$ is surjective. But, if $N$ is as above any nonzero element in $(A_{V})_{(\tilde{f}_{\sigma})}$ can be written in the form $\frac{a}{(\tilde{f}_{\sigma})^{kN}}$ for some $k>0$, where $a\in A$ is homogeneous of degree $kN\deg(\tilde{f}_{\sigma})$. Then $a$ belongs to the image of ${\mathcal{O}}(\widetilde{V})\otimes A$, which implies the desired surjectivity.